40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1079     Entry time: Thu Oct 23 21:52:27 2008
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: FSS UGF now 450kHz 
I measured the open loop transfer function of the FSS, for the first time after I mitigated the oscillation.
The attached plot shows the comparison of the OPLTF before and after the oscillation was mitigated.
Blue curves are when AD797 was oscillating, and the red ones are after AD797s were replaced by AD829s.
The FSS gain slider values are the same for the both measurements.
There is a notable difference in the shape of the TF.
Right now the UGF is around 450kHz with the phase margin of 50deg.
When the gain is increased by a few dBs in the common gain slider, the PC path becomes saturated.
This might be caused by the peak in the OPLTF at 1.7MHz sticking out of the 0dB line.
Another peak at 770kHz is also annoying.
Too bad that I did not take the TF above 1MHz before the oscillation was mitigated.
Also at 100kHz, the new TF has a lower gain than the old one, although it looks like the slope of the red curve is getting steeper and
it is catching up the blue one at lower frequencies.
I will measure the TF below 100kHz later.

With this bandwidth, I was able to increase the MC gain further.
I will report on the MC open loop measurements soon.
Attachment 1: FSS_OPLTF.png  80 kB  | Hide | Hide all
FSS_OPLTF.png
ELOG V3.1.3-